OpenStudy (anonymous):

Ali drinks three 12-ounce sodas every day and has done so for the past 2 years. His dentist warns him that acid and sugar from drinking that many sodas are damaging his teeth. Ali needs to reduce his intake to two 12-ounces sodas per week. By what percent does Ali need to reduce his soda intake for the week? A. 90 B. 67 C. 45 D. 80

OpenStudy (anonymous):

I would say A.

OpenStudy (anonymous):

Thank you very much for the help. Now I just have to figure out how you came up with the answer because I have to show my work.

OpenStudy (anonymous):

I did process of elimination so b and c are out. And since he's only drinking a 12 once soda a week it had to be 90 or a.

OpenStudy (anonymous):

Okay, thank you. That makes more sense to me now.

OpenStudy (anonymous):

Anytime! ☺

OpenStudy (a1234):

3 sodas every day means 21 sodas a week. His intake needs to be reduced to 2 a week. 2/21 = x/100 x is approximately 9.5. 100 - 9.5 = 90.5, A

OpenStudy (anonymous):

Thank you very much!